OpenStudy (shamil98):

I need some assistance with some calculus review problems. The function \[\ v(t) = t^2 - 8t + 15 \] is the velcotiy of a particle moving along the x - axis, where \[\ t \ge 0 \] and the velocity is measured in m/sec. (a) Determine when the particle is moving to the right, to the left, and stopped. (b) Find the particle's displacement for \[\ 0 \le t \le 6 \] Justify answers for (a) and (b). (c) Write but do not evaluate the integral used to find the total distance traveled by the particle for \[\ 0 \le t \le 6 \]

OpenStudy (shamil98):

I know how to do part c at the least... \[\text{Total Distance} = \int\limits_{0}^{6} | t^2 - 8t + 15| ~dt\]

OpenStudy (dan815):

okay the others are, its moving forward when velocity is positive and backward when its negative so

OpenStudy (dan815):

displacement u want to find the mean velocity and multiply by time

OpenStudy (dan815):

i think. id have to work it out to be sure

OpenStudy (anonymous):

Displacement can be computed by integrating \(v(t)\) over \(0\le t\le6\), since \(\int v(t)~dt=s(t)\) and \(s(t)\) is the displacement.

OpenStudy (shamil98):

Oh, i just plugged it in and added the values 3 + 15 = 18 gives the same value as I did when integrating it .

OpenStudy (shamil98):

so it's moving to the right when v(t) > 0 left when v(t) < 0 and stopped when v(t) = 0 so to the right would be: t > 5 , t < 3 to the left would be : 3 < t < 5 and stopped would be when t = 3 , 5

OpenStudy (shamil98):

is that correct ?

OpenStudy (anonymous):

yes

OpenStudy (shamil98):

alrighty, thanks both of you.